# Vertex F-algebras and their associated Lie algebra

### v1

@@ 3.3 Filtrations
A filtration on a ring $A$ is a descreasing sequence of subgroups
$\cdots\supset A_{(n)}\supset A_{(n+1)}\supset\cdots\qquad(n\in\mathbb{Z}),$ (3.7)
@@
Let $\mathfrak{a}\triangleleft A$ be an ideal. The $\mathfrak{a}$-adic filtration on $A$ is defined by $A_{(n)}=\mathfrak{a}^{n}$ for $n>0$ and $A_{(n)}=A$ for $n\leqslant 0.$ If $\mathfrak{a}=(0)$ this is the trivial filtration on $A,$ which induces the discrete topology.
For example, $A\llbracket z\rrbracket$ carries the $(z)$-adic filtration where $A\llbracket z\rrbracket_{(n)}=z^{n}A\llbracket z\rrbracket$ for $n\geqslant 0.$ We also call the filtration on $A\llparenthesis z\rrparenthesis$ defined by $A\llparenthesis z\rrparenthesis_{(n)}=z^{n}A\llbracket z\rrbracket$ for $n\in\mathbb{Z}$ the $(z)$-adic filtration, although $(z)$ is not an ideal in $A\llparenthesis z\rrparenthesis.$
@@ 3.4 Substitutions
The following is the main technical result for defining substitutions. Parts (a)–(c) extend [2, III,§2.6] to the case where $A_{(0)}\neq A.$
@@
(a) Let $A$ be a filtered ring. Then
$A\llparenthesis z\rrparenthesis_{(n)}=\Bigl{\{}\sum\nolimits_{i\in\mathbb{Z}}a_{i}z^{i}\enskip\Big{|}\enskip a_{i}\in A_{(n-i)}\Bigr{\}},\quad n\in\mathbb{Z},$ (3.8)
defines a filtration on the Laurent ring $A\llparenthesis z\rrparenthesis.$ If $A$ is separated, then so is $A\llparenthesis z\rrparenthesis$ and the power series ring $A\llbracket z\rrbracket$ is a closed subspace with filtration
$A\llbracket z\rrbracket_{(n)}=A\llparenthesis z\rrparenthesis_{(n)}\cap A\llbracket z\rrbracket.$ (3.9)
  1. (b)
    If $A$ is complete, then so are $A\llparenthesis z\rrparenthesis$ and $A\llbracket z\rrbracket.$
  2. (c)
    If $A_{(0)}=A,$ then the polynomials $A[z]$ are dense in $A\llbracket z\rrbracket$ and the Laurent polynomials $A[z^{\pm 1}]$ are dense in $A\llparenthesis z\rrparenthesis.$
  3. (d)
    (Universal properties) For each filtered morphism $f\colon A\to B$ into a complete ring $B$ and each invertible element $b\in B_{(1)}$ there is a unique filtered morphism $\bar{f}\colon A\llparenthesis z\rrparenthesis\to B$ with $\bar{f}|_{A}=f$ and $\bar{f}(z)=b.$ Here $A\llparenthesis z\rrparenthesis$ is given the filtration generated by (3.8) and the $(z)$-adic filtration. Pictorially,
    ${A\llparenthesis z\rrparenthesis}$${B.}$${A}$$\scriptstyle{z\mapsto b}$$\scriptstyle{f}$ (3.10)
    Similarly for multivariable Laurent rings $A\llparenthesis z_{1},\ldots,z_{n}\rrparenthesis$ and also for power series rings, where we drop the invertibility condition.
@@
(a) It is easy to check that (3.8) defines a filtration of $A\llparenthesis z\rrparenthesis,$ which is separated if the original filtration is so. Moreover, $A\llbracket z\rrbracket$ is a closed subspace.
(b) To prove that $A\llparenthesis z\rrparenthesis$ is complete, let $f_{n}(z)=\sum_{i\in\mathbb{Z}}a_{i,n}z^{i},n\in\mathbb{N},$ be a Cauchy sequence. For each $N$ there is $n_{0}$ such that for all $n,m\geqslant n_{0}$ and $i$ we have
$\forall n,m\geqslant n_{0}:a_{i,n}-a_{i,m}\in A_{(N-i)}.$ (3.11)
@@
$A[z]\longrightarrow B,\quad z\longmapsto b.$ (3.12)
First, put the trivial filtration on $A.$ Then (3.12) is filtered for the $(z)$-adic topology since $b\in B_{(1)}.$ As $B$ is complete, we may extend to a filtered map $\bar{f}$ on the completion $A\llbracket z\rrbracket.$ Observe that $\bar{f}$ is also automatically filtered for (3.8). Hence $\bar{f}$ is filtered for the generated filtration. Conversely, if $\bar{f}$ is filtered for the generated filtration, then it is also filtered for the $(z)$-adic filtration for which $A[z]$ is dense. This proves uniqueness. Similarly for Laurent series. ∎
@@
Applying (3.8) inductively starting with the trivial filtration on $R,$ we obtain the complete filtration on iterate Laurent rings $R\llparenthesis x_{1}\rrparenthesis\cdots\llparenthesis x_{m}\rrparenthesis,$ where to simplify notation we work in the single variable case. For $h(x_{1},\ldots,x_{m})$ to be in positive complete filtration means that each monomial occurring in $h$ has strictly positive total degree. We also have the generated filtration on $R\llparenthesis x_{1}\rrparenthesis\cdots\llparenthesis x_{m}\rrparenthesis,$ obtained by using the generated filtration inductively at each step.
@@
Let $f_{i}(y_{1},\ldots,y_{n})\in R\llparenthesis y_{1}\rrparenthesis\cdots\llparenthesis y_{n}\rrparenthesis,$ $1\leqslant i\leqslant m,$ be invertible and in positive complete filtration. There is a unique substitution morphism
$|_{x_{i}^{\pm 1}\to f_{i}(y_{1},\ldots,y_{n})^{\pm 1}}\colon R\llparenthesis x_{1}\rrparenthesis\cdots\llparenthesis x_{m}\rrparenthesis\longrightarrow R\llparenthesis y_{1}\rrparenthesis\cdots\llparenthesis y_{n}\rrparenthesis,\quad x_{i}\longmapsto f_{i}(y_{1},\ldots,y_{n}),$
@@
Moreover, given $g_{j}(z_{1},\ldots,z_{p})\in R\llparenthesis z_{1}\rrparenthesis\cdots\llparenthesis z_{p}\rrparenthesis,$ $1\leqslant j\leqslant n,$ invertible and in positive complete filtration we have functoriality
$|_{y_{j}^{\pm 1}\to g_{i}(z_{1},\ldots,y_{p})^{\pm 1}}\circ|_{x_{i}^{\pm 1}\to f_{i}(y_{1},\ldots,y_{n})^{\pm 1}}=|_{x_{i}^{\pm 1}\to h_{i}(z_{1},\ldots,z_{p})^{\pm 1}}$
where $h_{i}(z_{1},\ldots,z_{p})=f_{i}(y_{1},\ldots,y_{n})|_{y_{j}^{\pm 1}\to g_{i}(z_{1},\ldots,y_{p})^{\pm 1}}.$
@@
$R\llparenthesis v\rrparenthesis\llparenthesis w\rrparenthesis$ $\longrightarrow R\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis,$ $v$ $\longmapsto i_{z,w}F(z,w),$ (3.13)
$R\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis$ $\longrightarrow R\llparenthesis v\rrparenthesis\llparenthesis w\rrparenthesis,$ $z$ $\longmapsto i_{v,w}F(v,\iota(w)),$ (3.14)
@@
$R\llparenthesis v\rrparenthesis\longrightarrow R\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis,\quad v\longmapsto i_{z,w}F(z,w).$ (3.15)
@@
For vertex algebras one may restrict to substitutions in Laurent polynomials. As a formal group law may involve infinitely many powers, this is no longer possible for vertex $F$-algebras.
3.5 F-binomial coefficients
Let $F(z,w)$ be a formal group law. The lowest coefficients of $F(z,w),$ $F(z,\iota w),$ $F(\iota z,w),$ $\iota F(z,w)$ viewed in $R\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis$ are the units $z,\iota(z)\in R\llparenthesis z\rrparenthesis^{\times}.$ Hence they are invertible in $R\llparenthesis z\rrparenthesis\llparenthesis w\rrparenthesis$ and there are well-defined integer powers
$i_{z,w}F(z,w)^{n},\enskip i_{z,w}F(\iota z,w)^{n},\enskip i_{z,w}F(z,\iota w)^{n},\enskip i_{z,w}(\iota F(z,w))^{n},\quad\forall n\in\mathbb{Z}.$
@@ 3.5 F-binomial coefficients
$\binom{n}{i,j}_{\!\!F}$ $=0,$ if $j<0$ or $i+j<n,$ (3.17)
$\binom{n}{i,0}_{\!\!F}$ $=\delta_{i}^{m},$ (3.18)
$\binom{n}{i,j}_{\!\!F}$ $=\binom{n}{j,i}_{\!\!F},$ if $n\geqslant 0,$ (3.19)
$\binom{m+n}{r,s}_{\!\!F}$ $=\sum_{\begin{subarray}{c}i+k=r\\ j+\ell=s\end{subarray}}\binom{m}{i,j}_{\!\!F}\binom{n}{k,\ell}_{\!\!F}.$ (3.20)
@@
In particular, for the additive group law ($s=0$) we recover the ordinary binomial coefficients $\binom{n}{i,j}_{\!\!F_{a}}=\binom{n}{j}$ if $i+j=n$ and $\binom{n}{i,j}_{\!\!F_{a}}=0$ else. For the multiplicative group law, put $s=1.$
3.6 F-delta distributions
@@
$z^{-1}\delta_{F_{a}}\left(\frac{w}{z}\right)=i_{z,\underline{w}}(z-w)^{-1}-i_{w,\underline{z}}(z-w)^{-1}=\sum\nolimits_{n\in\mathbb{Z}}w^{n}z^{-n-1}.$ (3.21)
More generally, for the one-parameter formal group law (2.2) we have
@@
(a) The F-delta distribution is supported on the diagonal,
$\delta_{F}\left(\frac{w}{z}\right)f(z)=\delta_{F}\left(\frac{w}{z}\right)f(w),\qquad\forall f\in R\llparenthesis z\rrparenthesis.$ (3.22)
More generally, if $f(z,w)\in R\llparenthesis z,w\rrparenthesis$ converges on the diagonal,
@@
    $i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)-i_{z_{2},\underline{z}_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)\\ =i_{z_{1},\underline{z}_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)$
  2. (c)
@@
(a) The identities (3.22) and (3.23) are well-known for the classical delta distribution. By (2.10) and (3.6) the delta distributions are related by
$\delta_{F}\left(\frac{w}{z}\right)=G(z,w)\delta_{F_{a}}\left(\frac{w}{z}\right).$ (3.24)
Therefore (3.22) and (3.23) follow directly from the classical case.
@@
$i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)=z_{0}^{-1}\delta_{F}\left(\frac{w}{z_{0}}\right)\Big{|}_{w^{\pm 1}\to i_{z_{1},\underline{z}_{2}}F(z_{1},\iota\underline{z}_{2})^{\pm 1}}$
@@
$i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)=i_{z_{0},\underline{w}}F(z_{0},\iota\underline{w})^{-1}\big{|}_{\underline{w}\to i_{\underline{z}_{1},\underline{z}_{2}}F(\underline{z}_{1},\iota\underline{z}_{2})}\\ \hskip 85.35826pt-i_{w,\underline{z}_{0}}F(\underline{z}_{0},\iota w)^{-1}\big{|}_{w^{\pm 1}\to i_{z_{1},\underline{z}_{2}}F(z_{1},\iota\underline{z}_{2})^{\pm 1}}\\ =i_{z_{0},(\underline{z}_{1},\underline{z}_{2})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{1},(\underline{z}_{2},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1}.$
@@
$i_{z_{2},\underline{z}_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)$ $=i_{z_{0},(\underline{z}_{2},\underline{z}_{1})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{2},(\underline{z}_{1},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1}$
$i_{z_{1},\underline{z}_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)$ $=i_{z_{2},(\underline{z}_{1},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{1},(\underline{z}_{0},\underline{z}_{2})}F(z_{0},\iota z_{1},z_{2})^{-1}$
@@
Finally, using (3.22) put $z_{1}=F(z_{0},z_{2})$ into $\frac{F(\iota z_{0},z_{1},\iota z_{2})}{F(z_{0},\iota z_{1},z_{2})}$ on the left to get $1$
By (3.23) we may put $z=w$ in (3.24) and use (2.11) to get

### v2

@@ 3.3 Filtrations
A filtration on a ring $A$ is a decreasing sequence of subgroups
$\cdots\supset A_{(n)}\supset A_{(n+1)}\supset\cdots,\qquad n\in\mathbb{Z},$ (3.5)
@@
Let $\mathfrak{a}\triangleleft A$ be an ideal. The $\mathfrak{a}$-adic filtration on $A$ is defined by $A_{(n)}=\mathfrak{a}^{n}$ for n>0{n>0} and $A_{(n)}=A$ for $n\leqslant 0$. If $\mathfrak{a}=(0)$, this is the trivial filtration on $A$, which induces the discrete topology.
For example, $A\llbracket z\rrbracket$ carries the $(z)$-adic filtration where $A\llbracket z\rrbracket_{(n)}=z^{n}A\llbracket z\rrbracket$ for $n\geqslant 0$. We also call the filtration on A((z))A(\!(z)\!) defined by $A(\!(z)\!)_{(n)}=z^{n}A\llbracket z\rrbracket$ for $n\in\mathbb{Z}$ the $(z)$-adic filtration, although $(z)$ is not an ideal in A((z))A(\!(z)\!).
@@ 3.4 Substitutions
The following is the main technical result for defining substitutions. Parts (a)–(c) extend [2, Chapter III, Section 2.6] to the case where $A_{(0)}\neq A$.
@@
  • $(a)$
    Let $A$ be a filtered ring. Then
    $A(\!(z)\!)_{(n)}=\Bigl\{\sum_{i\in\mathbb{Z}}a_{i}z^{i}\,\Big|\,a_{i}\in A_{(n-i)}\Bigr\},\qquad n\in\mathbb{Z},$ (3.6)
    defines a filtration on the Laurent ring A((z))A(\!(z)\!). If $A$ is separated, then so is A((z))A(\!(z)\!) and the power series ring $A\llbracket z\rrbracket$ is a closed subspace with filtration
    $A\llbracket z\rrbracket_{(n)}=A(\!(z)\!)_{(n)}\cap A\llbracket z\rrbracket.$ (3.7)
  • $(2)$
    If $A$ is complete, then so are A((z))A(\!(z)\!) and $A\llbracket z\rrbracket$.
  • $(3)$
    If $A_{(0)}=A$, then the polynomials $A[z]$ are dense in $A\llbracket z\rrbracket$ and the Laurent polynomials $A\big[z^{\pm 1}\big]$ are dense in A((z))A(\!(z)\!).
  • $(4)$
    (Universal properties) For each filtered morphism $f\colon A\to B$ into a complete ring $B$ and each invertible element $b\in B_{(1)}$, there is a unique filtered morphism $\bar{f}\colon A(\!(z)\!)\to B$ with $\bar{f}|_{A}=f$} and $\bar{f}(z)=b$. Here A((z))A(\!(z)\!) is given the filtration generated by (3.6) and the $(z)$-adic filtration. Pictorially,
[…]
    Similarly for multivariable Laurent rings $A(\!(z_{1},\dots,z_{n})\!)$ and also for power series rings, where we drop the invertibility condition.
@@
(a) It is easy to check that (3.6) defines a filtration of A((z))A(\!(z)\!), which is separated if the original filtration is so. Moreover, $A\llbracket z\rrbracket$ is a closed subspace.
(b) To prove that A((z))A(\!(z)\!) is complete, let $f_{n}(z)=\sum_{i\in\mathbb{Z}}a_{i,n}z^{i}$, $n\in\mathbb{N}$, be a Cauchy sequence. For each $N$, there is $n_{0}$ such that for all $n,m\geqslant n_{0}$ and $i$ we have
$\forall n,m\geqslant n_{0}\colon\ a_{i,n}-a_{i,m}\in A_{(N-i)}.$ (3.9)
@@
$A[z]\longrightarrow B,\qquad z\longmapsto b.$ (3.10)
First, put the trivial filtration on $A$. Then (3.10) is filtered for the $(z)$-adic topology since $b\in B_{(1)}$}. Since $B$ is complete, we may extend to a filtered map $\bar{f}$ on the completion $A\llbracket z\rrbracket$. Observe that $\bar{f}$ is also automatically filtered for (3.6). Hence $\bar{f}$ is filtered for the generated filtration. Conversely, if $\bar{f}$ is filtered for the generated filtration, then it is also filtered for the $(z)$-adic filtration for which $A[z]$ is dense. This proves uniqueness. Similarly for Laurent series.∎
@@
Applying (3.6) inductively starting with the trivial filtration on $R$, we obtain the complete filtration on iterate Laurent rings $R(\!(x_{1})\!)\cdots(\!(x_{m})\!)$, where to simplify notation we work in the single variable case. For $h(x_{1},\dots,x_{m})$, to be in positive complete filtration means that each monomial occurring in $h$ has strictly positive total degree. We also have the generated filtration on $R(\!(x_{1})\!)\cdots(\!(x_{m})\!)$, obtained by using the generated filtration inductively at each step.
@@
Let $f_{i}(y_{1},\dots,y_{n})\in R(\!(y_{1})\!)\cdots(\!(y_{n})\!)$, $1\leqslant i\leqslant m$, be invertible and in positive complete filtration. There is a unique substitution morphism
$|_{x_{i}^{\pm 1}\to f_{i}(y_{1},\dots,y_{n})^{\pm 1}}\colon\ R(\!(x_{1})\!)\cdots(\!(x_{m})\!)\longrightarrow R(\!(y_{1})\!)\cdots(\!(y_{n})\!),x_{i}\longmapsto f_{i}(y_{1},\dots,y_{n}),$
@@
Moreover, given $g_{j}(z_{1},\dots,z_{p})\in R(\!(z_{1})\!)\cdots(\!(z_{p})\!)$, $1\leqslant j\leqslant n$, invertible and in positive complete filtration we have functoriality
$|_{y_{j}^{\pm 1}\to g_{i}(z_{1},\dots,y_{p})^{\pm 1}}\circ|_{x_{i}^{\pm 1}\to f_{i}(y_{1},\dots,y_{n})^{\pm 1}}=|_{x_{i}^{\pm 1}\to h_{i}(z_{1},\dots,z_{p})^{\pm 1}},$
where $h_{i}(z_{1},\dots,z_{p})=f_{i}(y_{1},\dots,y_{n})|_{y_{j}^{\pm 1}\to g_{i}(z_{1},\dots,y_{p})^{\pm 1}}$.
@@
$R(\!(v)\!)(\!(w)\!)\longrightarrow R(\!(z)\!)(\!(w)\!),\qquad v\longmapsto i_{z,w}F(z,w),$ (3.11)
$R(\!(z)\!)(\!(w)\!)\longrightarrow R(\!(v)\!)(\!(w)\!),\qquad z\longmapsto i_{v,w}F(v,\iota(w)).$ (3.12)
@@
$R(\!(v)\!)\longrightarrow R(\!(z)\!)(\!(w)\!),\qquad v\longmapsto i_{z,w}F(z,w).$ (3.13)
@@
For vertex algebras, one may restrict to substitutions in Laurent polynomials. As a formal group law may involve infinitely many powers, this is no longer possible for vertex $F$-algebras.
3.5 $\boldsymbol{F}$-binomial coefficients
Let $F(z,w)$ be a formal group law. As observed in Example 3.4, for all $n\in\mathbb{Z}$ there are well-defined integer powers
$i_{z,w}F(z,w)^{n},\qquad i_{z,w}F(\iota z,w)^{n},\qquad i_{z,w}F(z,\iota w)^{n},\qquad i_{z,w}(\iota F(z,w))^{n}.$
@@ 3.5 $\boldsymbol{F}$-binomial coefficients
$\binom{n}{i,j}_{\!\!F}=0\qquad\text{if}\ j<0\ \text{or}\ i+j<n,$
$\binom{n}{i,0}_{\!\!F}=\delta_{i}^{m},\qquad\binom{n}{i,j}_{\!\!F}=\binom{n}{j,i}_{\!\!F}\qquad\text{if}\ n\geqslant 0,$
$\binom{m+n}{r,s}_{\!\!F}=\sum_{\begin{subarray}{c}i+k=r\\ j+\ell=s\end{subarray}}\binom{m}{i,j}_{\!\!F}\binom{n}{k,\ell}_{\!\!F}.$
@@
In particular, for the additive group law ($s=0$), we recover the ordinary binomial coefficients $\binom{n}{i,j}_{\!\!F_{a}}=\binom{n}{j}$ if $i+j=n$ and $\binom{n}{i,j}_{\!\!F_{a}}=0$ else. For the multiplicative group law, put $s=1$.
3.6 $\boldsymbol{F}$-delta distributions
@@
$z^{-1}\delta_{F_{a}}\left(\frac{w}{z}\right)=i_{z,\underline{w}}(z-w)^{-1}-i_{w,\underline{z}}(z-w)^{-1}=\sum_{n\in\mathbb{Z}}w^{n}z^{-n-1}.$ (3.15)
More generally, for the one-parameter formal group law (2.2), we have
@@
  • $(a)$
    The F-delta distribution is supported on the diagonal,
    $\delta_{F}\left(\frac{w}{z}\right)f(z)=\delta_{F}\left(\frac{w}{z}\right)f(w),\qquad\forall f\in R(\!(z)\!).$ (3.16)
    More generally, if $f(z,w)\in R(\!(z,w)\!)$ converges on the diagonal,
@@
    $i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)-i_{z_{2},\underline{z}_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)=i_{z_{1},\underline{z}_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right).$
  • $(3)$
@@
(a) The identities (3.16) and (3.17) are well-known for the classical delta distribution. By (2.9) and (3.4), the delta distributions are related by
$\delta_{F}\left(\frac{w}{z}\right)=G(z,w)^{-1}\delta_{F_{a}}\left(\frac{w}{z}\right).$ (3.18)
Therefore, (3.16) and (3.17) follow directly from the classical case.
@@
$i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)=z_{0}^{-1}\delta_{F}\left(\frac{w}{z_{0}}\right)\Big|_{w^{\pm 1}\to i_{z_{1},\underline{z}_{2}}F(z_{1},\iota\underline{z}_{2})^{\pm 1}}.$
@@
$i_{z_{1},\underline{z}_{2}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)={}$ $i_{z_{0},\underline{w}}F(z_{0},\iota\underline{w})^{-1}\big|_{\underline{w}\to i_{\underline{z}_{1},\underline{z}_{2}}F(\underline{z}_{1},\iota\underline{z}_{2})}$
$-i_{w,\underline{z}_{0}}F(\underline{z}_{0},\iota w)^{-1}\big|_{w^{\pm 1}\to i_{z_{1},\underline{z}_{2}}F(z_{1},\iota\underline{z}_{2})^{\pm 1}}$
$={}$ $i_{z_{0},(\underline{z}_{1},\underline{z}_{2})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{1},(\underline{z}_{2},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1}.$
@@
$i_{z_{2},\underline{z}_{1}}z_{0}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{2})}{z_{0}}\right)$ $=i_{z_{0},(\underline{z}_{2},\underline{z}_{1})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{2},(\underline{z}_{1},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1},$
$i_{z_{1},\underline{z}_{0}}z_{2}^{-1}\delta_{F}\left(\frac{F(z_{1},\iota z_{0})}{z_{2}}\right)$ $=i_{z_{2},(\underline{z}_{1},\underline{z}_{0})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{1},(\underline{z}_{0},\underline{z}_{2})}F(z_{0},\iota z_{1},z_{2})^{-1}.$
@@
$i_{z_{2},\underline{z}_{0}}z_{1}^{-1}\delta_{F}\left(\frac{F(z_{0},z_{2})}{z_{1}}\right)\cdot\frac{F(\iota z_{0},z_{1},\iota z_{2})}{F(z_{0},\iota z_{1},z_{2})}$
$\qquad=i_{z_{1},(\underline{z}_{0},\underline{z}_{2})}F(z_{0},\iota z_{1},z_{2})^{-1}-i_{z_{2},(\underline{z}_{0},\underline{z}_{1})}F(z_{0},\iota z_{1},z_{2})^{-1}$
@@
Finally, using (3.16) put $z_{1}=F(z_{0},z_{2})$ into $\frac{F(\iota z_{0},z_{1},\iota z_{2})}{F(z_{0},\iota z_{1},z_{2})}$ on the left to get $1$. ∎
By (3.17), we may put $z=w$ in (3.18) and use (2.10) to get